# PAL: Probing Audio Encoders via LLMs - Audio Information Transfer into LLMs

### v2

@@ 3.3 LAL: Lightweight Audio-LLM Integration
Qwen-Audio-Chat-8B (Chu et al., 2023b) 3.5 3.3 3.6 3.5
@@
3.4 PAL: An Encoder Aware Architecture Extending LAL with Speech Understanding
Building on the LAL analysis and results, we ask when PLITS should be preferred over LAL. For the Whisper speech encoder (Radford et al., 2023), our initial experiments on emotion recognition and gender classification indicate that Whisper benefits from decoding inside the LLM; see Table 5. This aligns with classical neuro linguistics where Wernicke’s area is primarily involved in comprehension and has long been associated with processing language in both written and spoken forms, while the angular gyrus supports association across auditory, visual, and other sensory inputs. By analogy, speech features may become most useful when interpreted within a language context, whereas general audio benefits from a modality specific pathway.
Motivated by this, we introduce PAL (Probing the Audio Encoders via LLM), an encoder aware hybrid that chooses the integration per encoder. General audio encoders, SSLAM and CLAP, use LAL integration. The speech encoder, Whisper, uses PLITS integration (refer to Fig. 2 (C)).
Table 5: Integration choices for Whisper evaluated on IEMOCAP (Busso et al., 2008) (emotion recog.) and VoxCeleb2 (Hechmi et al., 2021) (gender cls.) (accuracy, %).
SSLAM+CLAP Integration Whisper Integration IEMOCAP Voxceleb2
PLITS PLITS 65.67 96.69
LAL LAL 66.88 97.19
LAL PLITS 68.81 97.99
Empirical Evaluation of PAL. We train PAL on a unified instruction tuning corpus covering speech, music, and general audio, and evaluate it on classification and reasoning benchmarks. As shown in Table 5 for classification and Tables 6 and 7 for reasoning, PAL is comparable to PLITS in accuracy while retaining efficiency advantages. We also observe that adding a Whisper encoder changes performance in the general audio(sound) and music domains. We hypothesize that this is because Whisper encodes background sounds, as reported by Gong et al. (2023a), which provides some event detection capability.
Our PAL versus PLITS comparison is controlled within our setup, using the same backbone, data, and training hyperparameters; see Appendix C.2 for details. The primary comparison in these tables is therefore between PAL and PLITS, and results from prior work are included only to place PAL in the broader literature. With the exception of Audio Flamingo 2, the other systems are based on PLITS. The higher scores reported by some prior systems over PLITS largely reflect larger training sets, larger LLMs, and stronger audio encoders. This work assesses the integration in isolation, which is why we focus on PAL versus PLITS comparison.
Table 6: Evaluation on MMAU-v05.15.25 (Sakshi et al., 2024) (accuracy, %). Sound (Sn), Music (Mu), Speech (Sp), and overall Average. Except for Audio Flamingo 2, all other systems use PLITS; their higher scores mainly stem from larger datasets, bigger LLMs, and stronger audio encoders. Boldface is used only for fair comparisons.
@@ 3.4 PAL: An Encoder Aware Architecture Extending LAL with Speech Understanding
Qwen2.5-Omni-7B (Xu et al., 2025a) 78.10 76.77 65.90 67.33 70.60 68.90 71.50 71.00
Qwen2-Audio-Instruct-7B (Chu et al., 2024a) 67.27 61.17 56.29 56.29 55.67 55.57 59.90 57.40
@@
PLITS-1B (Baseline) 71.17 72.20 71.56 69.66 53.45 54.31 65.40 64.61
PAL-1B (Ours) 72.07 70.63 70.66 66.10 53.45 53.28 65.40 63.45
Table 7: Evaluation of PAL on  MMAR (Ma et al., 2025) (accuracy, %). Abbr: Sound (Sn), Music (Mu), Speech (Sp), and Total. Except for Audio Flamingo 2, all other systems use PLITS; their higher scores mainly stem from larger datasets, bigger LLMs, and stronger audio encoders. Boldface is used only for fair comparisons.
@@
PLITS-1B (Baseline) 38.79 42.72 40.48 18.18 44.50 39.02 41.67 41.20
PAL-1B(Ours)  40.61 41.75 38.10 36.36 45.87 52.44 41.67 42.20
4 Experimental Setup
4.1 LAL: Experimental Setup
Training Protocol. We train the proposed audio LLM variants on the one of the largest general audio instruction tuning datasets OpenAQA dataset (Gong et al., 2024) and CompA-R Ghosh et al. (2024). Our two-stage pipeline comprises: (i) connector pretraining, where only the connector is trained and all other modules are frozen; and (ii) joint training of the connector and the LLM. The audio encoders remain frozen throughout.
For reasoning and open ended question answering we additionally train on open ended data from OpenAQA Gong et al. (2024) as Stage 3 and on the reasoning dataset CompA R Ghosh et al. (2024) as Stage 4. Additional training details are in Appendix C.1.
Evaluation Protocol. To assess how effectively LAL transfers critical audio-event information from the encoder to the LLM’s latent space, we evaluate on downstream classification, captioning, and reasoning tasks. Following the LTU framework (Gong et al., 2024): (i) for classification, we measure semantic similarity by encoding both model text outputs and target audio labels with gpt-text-embedding-ada; (ii) for captioning, we use standard audio captioning datasets and report CIDEr and SPICE.
For reasoning, we adopt the compA-R-test and the evaluation protocol of (Ghosh et al., 2024): we prompt a text-only GPT-4 judge with the audio-LLM’s output and auxiliary metadata about the audio events, and obtain scores for Helpfulness, Clarity, Correctness, Depth, and Engagement. Additional evaluation details are in Appendix D.1.
4.2 PAL: Experimental Setup
Training Protocol. PAL follows the same two stage procedure as LAL: (i) connector pretraining, where only the connector is trained and all other modules are frozen; and (ii) joint training of the connector and the LLM. The audio encoders remain frozen throughout. For Stage 1, we construct a mixture from the general audio OpenAQA Stage 1 set, augmented with the OpenASQA (Gong et al., 2023b) Stage 1 split for speech understanding. For Stage 2, we use a curated audio, speech, and music reasoning instruction tuning corpus, specifically a 6M subset of AudioSkills (Goel et al., 2025).
Evaluation Protocol. We first target speech understanding with two tasks: speech recognition and speaker gender classification (using gpt-text-embedding-ada as explained in Section 4.1); These are evaluated after Stage 1 to assess how well the newly added Whisper encoder integrates with the LLM. We then assess general audio, music, and speech reasoning on MMAR and MMAU, which report detailed category wise performance.
5 Conclusion
We introduce LAL, which injects audio only through attention keys and values and skips feedforward processing for audio tokens. This reduces attention interactions and activations, yielding up to 64.1% lower memory and up to 247.5% higher training throughput with comparable performance as PLITS, the SOTA baseline integration on classification, captioning, and reasoning. We also propose PAL, an encoder aware hybrid that uses LAL for SSLAM and CLAP and PLITS for Whisper as it benefits from the decoding inside the LLM. LAL is a core architectural change rather than a parameter efficient fine tuning method, so the efficiency gains hold at inference and during training. For future work, we plan to scale to larger backbones, use higher quality instruction data to improve reasoning, and explore streaming and long context audio.
Acknowledgments
This research was supported by the EPSRC and BBC Prosperity Partnership “AI4ME: Future Personalized Object Based Media Experiences Delivered at Scale Anywhere” (EP/V038087/1). For part of the experiments, we used resources provided by the EuroHPC Joint Undertaking, which granted this project access through a EuroHPC Development Access call to the LEONARDO EuroHPC supercomputer hosted by CINECA (Italy) and to the resources of the LEONARDO consortium.
Ethics Statement
All experiments use publicly available datasets. The proposed approach enables beneficial applications, but it could also be misused, for example to monitor individuals without consent. We acknowledge these risks and will release code and models with care, including clear documentation and use guidance to support responsible research.
Reproducibility Statement
Implementation details are provided in Sections 3.3 and 3.4. Training details appear in Appendix C, and the evaluation protocol is described in Appendix D. Code and pretrained models will be made available upon acceptance.
Appendix A Appendix
Appendix B LLM Usage
Large language models were used only as assistive tools for editing and polishing text. We followed the benchmark protocol of  Ghosh et al. (2024) to rate audio LLM outputs; the GPT based evaluation is part of that benchmark. See Section 4.1 for details. LLMs were not used for model design, data selection, experiment setup, implementation, analysis, or generation of results. All technical content was written and verified by the authors.
Appendix C Training Details
C.1 LAL Training Details
We use OpenAQA (Gong et al., 2024) two stage training setup for LAL to report the results in Table 1. We also train on broader open ended data from OpenAQA (Gong et al., 2024) and on the reasoning dataset CompA R (Ghosh et al., 2024), with evaluations shown in Table 2. Additional training hyperparameters appear in Table 8.
Table 8: Hyper-parameters used for the three stage training of LAL and PLITS (Llama3.2 1B)
Training Configuration Stage 1 Stage 2 Stage 3 $|$ Stage 4
(Connector Pre training) (LLM Fine tuning) (LLM Fine tuning)
Optimizer AdamW (Loshchilov & Hutter, 2017)
Learning Rate Schedule Cosine (Loshchilov & Hutter, 2016)
Peak Learning Rate 0.001 0.0001 0.0001
Epochs 1 1 1
Warm up Ratio (steps) 0.05 0.03 0.03
Dataset Size 1.2 M 1.9 M 5.6 M $|$ 200 K
Batch Size 3232 1212 1212
Gradient Accumulation Steps 4
GPUs $2\times$ Nvidia A100 (80GB)
RAM 150 GB
Loss Next token loss on text part
C.2 PAL Training Details
PAL uses a two stage training protocol(Table  9). In Stage 1, we start from the Stage 1 dataset used for LAL and augment it with additional speech focused data from OpenASQA (Gong et al., 2023b). In Stage 2, we fine tune on a curated audio, speech, and music reasoning instruction corpus, AudioSkills (Goel et al., 2025). We use a 6M example subset of AudioSkills (from the original 10M) due to the unavailability of original audio files for some source datasets.
Table 9: Hyperparameters used for the two stage training of PAL and PLITS (Llama3.2 1B)
Training Configuration Stage 1 Stage 2
(Connector Pre training) (LLM Fine tuning)
Optimizer AdamW (Loshchilov & Hutter, 2017)
Learning Rate Schedule Cosine (Loshchilov & Hutter, 2016)
Peak Learning Rate 0.001 0.0001
Epochs 1 1
Warm up Ratio (steps) 0.05 0.03
Dataset Size 1.7 M 6.4 M
Batch Size 1616 44
Gradient Accumulation Steps 22 3232
GPUs $4\times$ Nvidia A100 (64GB)
RAM 250 GB
Loss Next token loss on text part
Appendix D Evaluation Details
D.1 LAL Evaluation Details
We follow the evaluation protocol of Gong et al. (2024) for classification and captioning, and use the CompA R test set of Ghosh et al. (2024) for reasoning. Below we summarize the datasets included in the Gong et al. (2024) protocol.
D.2 PAL Evaluation Details
For speech classification (emotion recognition and gender classification), we follow the protocol of Gong et al. (2023b). For combined sound, speech, and music reasoning, we evaluate on the standard benchmark datasets MMAU (Sakshi et al., 2024) and MMAR (Ma et al., 2025).
Appendix E LAL Integration with Frozen LLM FFN
Table 10: Performance evaluation of the LAL Integration with frozen FFN. Evaluation follows the protocol of Gong et al. (2024). AC: Audio caps, CL:Clotho AS2M: AudioSet 2M indicates CIDEr and indicates SPICE. Metrics: accuracy (ESC-50, VocalSound), Mi-F1 (DCASE), and mAP (FSD, AudioSet). Complete evaluation methodology explained in Section 4.1 and dataset details in Appendix D
LLM FFN PLITS LAL LFST Classification Captioning
Backbone Frozen ESC50 DCASE VS FSD AS2M AC CL AC CL
Llama3.2-1B 64.45 37.69 51.57 25.23 9.08 0.59 0.34 16.30 10.96
 76.70  40.97  60.87  31.44  11.83  0.66 0.38  16.97  11.87
71.80 33.99 55.28 29.38 10.48 0.63  0.40 16.11 11.75
E.1 LFST Connector: Language aligned and Fine grained Spatiotemporal Connector
We adopt the connector proposed in Cambrian (Tong et al., 2024) and apply it in our audio setting to fuse a language aligned encoder such as CLAP with a self supervised encoder such as SSLAM. The connector produces a compact set of latent tokens that combine semantic cues from CLAP with fine grained spatiotemporal detail from SSLAM, while keeping sequence length fixed and avoiding the overhead of naive concatenation.

### v3

@@ Appendix G Extending Table 1 and Table 2 with Prior audio-LLM performances
Qwen-Audio-Chat-8B (Chu et al., 2023) 3.5 3.3 3.6 3.5
@@
Appendix H Extending Table 3 and Table 4 with Prior audio-LLM performances
This section extends the MMAU, MMAR, and MMSU evaluations by comparing our PLITS, LAL, and PAL unified models to representative prior audio-LLMs, as reported in Tables 12, 13, and 14. We emphasize that these cross-paper comparisons are not strictly apples-to-apples: aside from Audio Flamingo 2, most prior systems rely on the standard PLITS integration, and their reported performance is often shaped by substantial differences in training data scale, instruction-tuning mixtures, LLM capacity, audio encoder strength etc. Accordingly, we include these results primarily to contextualize the absolute performance range of existing audio-LLMs on MMAU/MMAR/MMSU and to situate our methods within this broader landscape, rather than to claim a fully controlled architectural comparison against any single prior model.
Table 12: Evaluation on MMAU-v05.15.25 (Sakshi et al., 2024) (accuracy, %). Sound (Sn), Music (Mu), Speech (Sp). Note that values are missing (-) for LAL+ 3B as the MMAU-test online challenge was closed at the time of evaluation. Except for Audio Flamingo 2, all other systems use PLITS; their higher scores mainly stem from larger datasets, bigger LLMs. LT indicates the same model trained for more steps (no architectural changes); excluded from bolding to ensure a fair comparison.
@@ Appendix H Extending Table 3 and Table 4 with Prior audio-LLM performances
Qwen2.5-Omni-7B (Xu et al., 2025) 78.10 76.77 65.90 67.33 70.60 68.90 71.50 71.00
Qwen2-Audio-Instruct-7B (Chu et al., 2024) 67.27 61.17 56.29 56.29 55.67 55.57 59.90 57.40
@@
PLITS-3B 75.68 72.03 70.96 69.63 46.25 46.48 64.30 62.91
LAL-3B 73.57 69.53 67.66 69.67 47.15 50.07 62.80 63.24
LAL-3B (LT) 75.08 - 70.66 - 48.05 - 64.60 -
PAL-3B 78.08 74.60 73.05 71.30 59.46 58.36 70.20 68.20
Table 13: Evaluation of PAL on  MMAR (Ma et al., 2025) (accuracy, %). Abbr: Sound (Sn), Music (Mu), Speech (Sp). Except for Audio Flamingo 2, all other systems use PLITS; their higher scores mainly stem from larger datasets, bigger LLMs. LT indicates the same model trained for more steps (no architectural changes); excluded from bolding to ensure a fair comparison.
@@
PLITS-3B 38.79 40.29 37.41 36.36 48.17 40.24 50.00 41.10
LAL-3B 38.18 42.72 35.37 18.18 41.28 39.02 41.67 38.90
LAL-3B (LT) 47.88 43.69 38.10 27.27 46.44 45.12 54.17 43.50
PAL-3B 44.85 41.75 43.54 27.27 51.83 47.78 41.67 45.40
Table 14: Evaluation on MMSU (Wang et al., 2025) across perception and reasoning dimensions in Semantics (Seman.), Phonology (Phono.), and Paralinguistics (Para.) domains. Except for Audio Flamingo 2, all other systems adopt PLITS; most are trained on substantially larger, speech-heavy datasets than ours and use larger LLMs. LT indicates the same model trained for more steps (no architectural changes); excluded from bolding to ensure a fair comparison.
Perception Reasoning Avg
Model Seman. Phono. Para. Avg Seman. Phono. Para. Avg All
LTU-7B (Gong et al., 2024) 21.34 22.46 18.73 20.81 22.65 25.53 24.74 24.37 22.61
LTU-AS (Gong et al., 2023) 25.89 24.71 21.64 24.13 26.53 25.68 25.04 25.92 25.03
SALMONN-13B (Tang et al., 2024) 31.55 29.08 28.71 29.83 36.43 26.22 25.26 30.04 30.01
Qwen2-Audio-Instruct-7B (Chu et al., 2024) 52.14 32.87 35.56 39.02 77.62 64.81 46.67 68.90 53.27
Qwen2.5-Omni-3B (Xu et al., 2025) 52.04 38.73 39.19 42.37 81.20 81.12 41.19 72.76 56.83
Qwen2.5-Omni-7B (Xu et al., 2025) 55.12 37.33 39.35 42.50 88.00 81.37 48.36 79.83 60.57
Audio Flamingo-3-8.2B  (Goel et al., 2025) - - - - - - - - 61.4
PLITS-3B 24.09 31.98 34.75 31.12 49.46 52.3 42.99 49.71 40.12
LAL-3B 26.61 33.8 31.19 31.01 50.00 54.35 52.54 52.11 41.22
LAL-3B (LT) 31.34 34.33 27.13 30.78 53.97 58.75 51.34 55.54 42.76
PAL-3B 29.76 37.65 37.52 35.66 58.12 59.16 49.85 57.4 46.18
Appendix I Extended literature review
Audio-LLM Datasets
Beyond architecture, recent works have focused on high-quality instruction tuning datasets, both open-source and proprietary (Goel et al., 2025; Ghosh et al., 2024; Chu et al., 2024; Xu et al., 2025) and build audio reasoning benchmarks (Sakshi et al., 2024; Deshmukh et al., 2025a, b). Training PLITS or Flamingo-style models on these resources improves instruction following and audio reasoning, with most gains driven by the data rather than the integration scheme.
Audio Representation learning
To obtain rich semantic audio representations, recent advances in audio representation learning have led to powerful audio encoders trained with diverse objectives across different pretraining paradigms. The studies have shifted from simple supervised learning paradigms  (Gong et al., 2022a, 2021) to more complex self-supervised paradigms  (Huang et al., 2022; Ahmed et al., 2024; Chen et al., 2024; Alex et al., 2025) that employ contrastive objectives and masked-token prediction strategies to capture both global semantic structure and fine-grained local details within audio representations. Furthermore, in the multimodal pretraining paradigm, language-aligned audio representations are obtained through contrastive audio–language models (Elizalde et al., 2023; Wu et al., 2023; Ghosh et al., 2025b) which align the representations of audio and language into a unified semantic space. Transcription-based approaches  (Radford et al., 2023) leverage next token prediction on speech-to-text tasks to learn robust audio representations that capture speech semantics and acoustic-linguistic relationships.
Appendix J LFST Connector: Language aligned and Fine grained Spatiotemporal Connector
We adopt the connector proposed in Cambrian (Tong et al., 2024) and apply it in our audio setting to fuse a language aligned encoder such as CLAP (Elizalde et al., 2023) with a self supervised encoder such as SSLAM (Alex et al., 2025). The connector produces a compact set of latent tokens that combine semantic cues from CLAP with fine grained spatiotemporal detail from SSLAM, while keeping sequence length fixed and avoiding the overhead of naive concatenation.